# GWQ: Gradient-Aware Weight Quantization for Large Language Models

### v2

@@ 4 Experiments
Original 16 - 5.47 6.97 4.88 6.47 6.23 8.99
+ GPTQ 4 1024 5.83 7.79 5.12 7.72 8.21 11.48
+ AWQ 4 512 5.60 7.70 4.97 7.10 6.64 10.69
+ SPQR 4.63 1024 5.53 7.02 4.93 6.50 6.42 9.23
+ SPQR 3.98 1024 5.55 7.06 4.95 6.53 6.45 9.26
+ GWQ-O (Ours) 4.63 1 5.51 7.01 4.91 6.48 6.40 9.21
+ GWQ-R (Ours) 3.98 1 5.53 7.04 4.94 6.52 6.43 9.25
Table 3: Performance of the Llama-2-hf family and Llama-3-8B-hf on zero-shot tasks on language task datasets. GWQ demonstrates a lower loss of accuracy compared to other quantization methods.
Llama-2-7B-hf Llama-2-13B-hf Llama-3-8B-hf
Acc.(%$\uparrow$ Avg_bit cab_num WikiText2 C4 WikiText2 C4 WikiText2 C4
Original 16 - 61.01 56.77 62.88 57.90 58.34 51.71
+ GPTQ 4 1024 59.88 54.74 61.91 56.18 43.78 47.75
+ AWQ 4 512 60.54 55.16 62.53 56.44 57.07 48.96
+ SPQR 4.63 1024 60.80 56.63 62.71 57.78 57.70 51.20
+ SPQR 3.98 1024 60.56 55.18 62.55 56.45 57.09 48.97
+ GWQ-O (Ours) 4.63 1 60.83 56.65 62.73 57.80 57.71 51.22
+ GWQ-R (Ours) 3.98 1 60.58 55.20 62.56 56.47 57.11 48.99
Overview. First, we validate the superiority of our method over other quantization approaches on the Llama-2 family (Touvron et al., 2023b), providing detailed compression metrics for comparison. Additionally, we present the results of our method compressed to 2 bits versus 3 bits, confirming its reliability at lower bit precision. Following this, we demonstrate the quantization performance of our method on other models, showcasing its generalizability. Finally, we present zero-shot quantization results on multimodal models to verify the multimodal context generalizability of our approach.
4.1 Experiments Setting
Setups. The language-only models we selected include Llama2-7B and 13B (Touvron et al., 2023b), Falcon-7B (Penedo et al., 2023), and Mistral-7B-v0.1 (Jiang et al., 2023). These models were quantized using the first sample of the RedPajama dataset. To evaluate Perplexity (PPL) and Accuracy (Acc.), we utilize the WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020) validation sets. All experiments were conducted on an NVIDIA A100 (80G) GPU.
Baselines. Our primary baselines include AWQ (Lin et al., 2024) and GPTQ (Frantar et al., 2022a), while we benchmark our approach against the state-of-the-arts SPQR (Dettmers et al., 2023). To ensure that the average bit-width of the model remains below 4 bits—surpassing GPTQ and AWQ in compression—we quantize the non-outlier weights at 4-bit and 3-bit precision. These configurations are denoted GWQ-O for 4-bit quantization, GWQ-R for 3-bit quantization.
@@ 4.2 Main Result
Table 4: Perplexity of Qwen-VL family models on language tasks datasets. GWQ quantized model has a lower PPL compared to other current quantization methods, and the PPL also remains essentially constant at lower bits.
QWEN-VL QWEN-VL-chat
[…]
Original 16 8.21 10.10 9.85 11.87
+ GPTQ 4 9.31 11.97 11.50 13.59
+ AWQ 4 8.69 10.98 11.02 12.88
+ SPQR 4.63 8.54 10.25 10.13 12.02
+ GWQ-O (Ours) 4.63 8.24 10.10 9.86 11.88
+ GWQ-R (Ours) 3.63 8.53 10.42 10.94 12.62
Table 5: Performance of the Qwen-VL family on the zero-shot tasks on language tasks datasets. GWQ has lower loss of accuracy compared to other quantization methods.
QWEN-VL QWEN-VL-chat
Acc.(%$\uparrow$ Avg_bit WikiText2 C4 WikiText2 C4
Original 16 54.18 49.51 52.06 48.07
+ GPTQ 4 52.07 47.08 49.83 46.09
+ AWQ 4 53.58 48.12 50.77 47.43
+ SPQR 4.63 54.05 49.26 51.99 47.35
+ GWQ-O (Ours) 4.63 54.06 49.50 52.03 48.07
+ GWQ-R (Ours) 3.63 53.87 49.32 52.01 49.98
Table 6: Performance of the Qwen-VL family on the zero-shot tasks on target detection datasets. In the three RefCOCO datasets, GWQ is more accurate compared to SPQR for the same bits, achieving less quantization loss.
QWEN-VL QWEN-VL-chat
Acc.(%$\uparrow$ Avg_bit test-A test-B val test-A test-B val
Original 16 92.26 85.34 89.36 92.27 84.51 88.55
+ GPTQ 4 90.61 83.14 87.13 90.33 82.27 86.67
+ AWQ 4 91.33 84.51 88.01 91.14 83.61 87.77
+ SPQR 4.63 92.17 84.87 88.97 91.76 84.20 88.48
+ SPQR 3.98 91.33 84.51 88.01 91.15 83.61 87.77
+ GWQ-O (Ours) 4.63 92.18 84.88 88.98 91.78 84.22 88.49
+ GWQ-R (Ours) 3.98 91.34 84.52 88.03 91.16 83.62 87.79
Table 7: Comparison of speedup and memory consumption of different quantization methods. GWQ significantly improves the inference speed and reduces the memory required compared to the original model.
Llama2-7B-hf Llama2-13B-hf
Methods Avg_bit Speedup($\uparrow$) Memory($\downarrow$) Speedup($\uparrow$) Memory($\downarrow$)
Original 16 $\times$1 12.8 $\times$1 23.6
+ GPTQ 4 $\times$2.0 3.1 $\times$2.0 6.7
+ AWQ 4 $\times$2.1 3.1 $\times$2.1 6.7
+ SPQR 4.63 $\times$1.2 N/A $\times$1.2 N/A
+ GWQ (Ours) 4.63 $\times$1.2 4.1 $\times$1.2 7.1
Language Model. As shown in Tab. 2 and Tab. 3, we first evaluate the perplexity of the quantized model on the WikiText2 Merity et al. (2016) and C4 Raffel et al. (2020) datasets on the Llama2-7B and 13B models. Additionally, we assess the models’ ability to zero-shot on these three datasets and compare the results to several baseline methods. GWQ achieves a lower average bit-width than other quantization methods by applying more aggressive compression to the non-outlier weights. Compared to other quantization approaches, GWQ yields a lower PPL on all datasets under identical bit compression. For the zero-shot task, the Acc. of the quantized model of GWQ is significantly higher than that of other quantization methods. indicating that GWQ results in less performance degradation. Notably, under extreme compression, GWQ sustains comparable performance to 4-bit quantization, even at 3-bit.
Table 8: Performance of GWQ at 4bit quantization for different number of calibration samples. The quantization loss of the model essentially does not vary with the number of different calibration samples, allowing for low-loss quantization with fewer samples.
num_sample=1 num_sample=4 num_sample=8 num_sample=12
[…]
Llama2-7B-hf 5.51 7.01 5.50 7.01 5.51 7.00 5.50 6.99
Llama2-13B-hf 4.91 6.48 4.90 6.49 4.91 6.47 4.91 6.48
Llama-3-8B-hf 6.40 9.21 6.39 9.20 6.38 9.20 6.40 9.20
Table 9: Performance of GWQ at 4bit quantization for different sample of calibration samples. The model quantization loss is essentially invariant to changes in the quantization sample and is less dependent on sample type.
sample one sample two sample three sample four
[…]
Llama2-7B-hf 5.51 7.01 4.91 6.97 5.47 6.97 5.47 6.97
Llama2-13B-hf 4.91 6.48 4.88 6.47 4.88 6.47 4.88 6.47
Llama-3-8B-hf 6.40 9.21 5.25 7.75 5.25 7.75 5.25 7.75
@@
Multimodal Models. For the multimodal model, we evaluate the performance of the compressed model in language tasks and multimodal tasks with the Qwen-VL family (Bai et al., 2023a, b). Similar to the language models, we evaluate the language generation capabilities of the quantized model on the WikiText2, C4 dataset. The multimodal tasks are mainly focused on zero-shot target detection. We choose the RefCOCO (Kazemzadeh et al., 2014; Yu et al., 2016) dataset to evaluate the zero-shot capability of the quantized model compared to other quantization methods. Intersection over Union $IoU$, which is denoted as:
@@
As it shown in Tab. 4 and Tab. 5, quantized qwen-vl family model of GWQ outperforms the current stat-of-the-arts in both the language task and the zero-shot test, showing better performance. The GWQ quantized MLLM performs better on the verbal task compared to other quantization methods, so we will use the quantized model to test it on the visual grounding task. As it shown in Tab. 6, in the zero-shot grounding detection task, both GWQ-quantified Qwen-VL family models with currently outperform the current stat-of-the-arts on most of the datasets, achieving a generalization of the multimodal model, The performance is shown in Fig. 3, for the task of visual grounding detection of long context, the GWQ post-quantization model is more advantageous than SPQR post-quantization.
@@
Model Efficiency Comparison. For the model throughput (tokens/s), the speedup rate can be obtained by calculating the time required to generate 512 tokens after quantization and comparing it with the original model. For the model inference overhead (G), we use the peak value when the model generates 512 tokens as the maximum inference memory by recording it.
As illustrated in Tab. 7, the GWQ quantized model can achieve about 1.2$\times$ speedup relative to the original model, which is on par with the current mixed-accuracy method SPQR. Compared to the full-precision quantization method, GWQ results in higher inference latency due to floating-point conversion during inference. Compared to the original model, GWQ consumes less memory during inference, which makes it easier to deploy on edge devices.
Figure 4: Effect of calibration sample on quantification. For different numbers of calibration sets, the PPL of the quantized model tends to be essentially constant on WikiText2 and C4. Therefore, the effect of the number of calibration set on the quantization effect is negligible.
@@ 4.3 Ablation Study
In this section, we examine the impact of varying the number of calibration samples on the performance of our method. Tab. 8 presents the results for the calibration sample sizes 1, 4, 8 and 12, evaluated on the WikiText2 and C4 datasets. Tab. 8 and Fig. 4 indicate that increasing the number of calibration samples has minimal impact on the PPL of the quantized model. This suggests that the size of the calibration set has little influence on the quantization performance of GWQ. Compared to most of the current quantization methods, which require large datasets, GWQ has a lower calibration cost and can achieve calibration with fewer samples.
To further investigate the impact of different calibration samples, we selected four distinct samples, each of which was used as a calibration set to quantize the model. According to Tab. 9, when different calibration samples are used to quantize the same model, the resulting PPL values on the same dataset remain similar, indicating that the choice of calibration samples has a negligible effect on the quantization results. Unlike AWQ, GWQ can be well adapted to the calibration sets of different samples, and the quantized model performs well even if the calibration sets differ greatly.

### v3

@@ 4.1 Experiments Setting
$\text{PPL}=\exp\left(-\frac{1}{N}\sum_{i=1}^{N}\log P(x_{i}\mid x_{0:i-1})\right),$ (11)
where $x_{i}$ represents the predicted token conditioned on the former tokens ($x_{<i}$) in inference step $i$ and $c_{w}$ denotes the average characters per word in the generated sequence.
The next token prediction accuracy can be expressed as Eq. (18),
$\mathrm{Acc}=\frac{N_{\hat{y_{t}}=\overrightarrow{y_{t}}}}{T},$ (12)
where $\hat{y_{t}}$ is the predicted token by LLMs, $\overrightarrow{y_{t}}$ is the ground truth token, $T$ is the number of all tokens.
The exact calculation of next token prediction accuracy will be explained in A.1.
Baselines. Our baseline is categorized into uniform precision quantization and mixed precision quantization. For uniform precision quantization, we choose OPTQ Frantar et al. (2022b), QUIP Chee et al. (2023), GPTQ Frantar et al. (2022a) and AWQ Lin et al. (2024b) as the baseline. For mixed precision quantization, we choose SPQR Dettmers et al. (2023) as the baseline. For fairness, we test the results of SPQR and GWQ with an average bit less than 4 respectively.
@@ 4.2 Main Result
Table 4: Results of Llama family on massive multitask language understanding task. GWQ outperforms other quantization methods, especially when the average bit-width of the model is significantly lower.
Accuracy (%)
Model Methods
Average Bit
Calib. Num
Average STEM Humanities
Social Sciences
Other
Llama-2-7B-hf Original 16 - 45.96 37.04 43.38 51.84 52.44
+OPTQ 4 1024 33.97 25.34 32.41 38.89 39.22
+QuIP 4 1024 34.58 25.90 33.12 39.02 40.29
+GPTQ 4 1024 36.26 27.95 34.55 40.95 41.59
+AWQ 4 512 38.64 30.13 36.85 43.35 44.24
+SPQR-O 4.63 1024 41.34 33.63 39.38 46.65 47.25
+SPQR-R 3.98 1024 38.66 30.40 36.35 43.57 44.33
+GWQ-O 4.63 1 43.36 34.59 40.94 48.32 50.34
+GWQ-R 3.98 1 39.90 31.89 37.54 45.09 45.09
Llama-2-13B-hf Original 16 - 55.68 44.27 54.43 63.41 60.76
+OPTQ 4 1024 43.32 30.85 39.43 52.59 50.44
+QuIP 4 1024 53.54 31.59 40.22 53.90 51.44
+GPTQ 4 1024 45.47 33.45 41.49 54.46 52.49
+AWQ 4 512 47.35 35.48 43.46 56.1 54.33
+SPQR-O 4.63 1024 49.85 37.59 46.45 58.44 56.93
+SPQR-R 3.98 1024 47.90 35.94 44.23 56.57 54.87
+GWQ-O 4.63 1 52.35 40.82 49.22 61.07 59.34
+GWQ-R 3.98 1 49.04 36.95 45.95 57.35 55.90
Llama-2-13B-hf Original 16 - 65.01 55.39 56.57 76.87 71.35
+OPTQ 4 1024 51.90 43.59 44.49 61.59 57.93
+QuIP 4 1024 54.12 45.93 46.29 64.95 59.35
+GPTQ 4 1024 56.00 47.84 48.35 66.49 61.35
+AWQ 4 512 57.84 49.24 50.43 68.44 63.24
+SPQR-O 4.63 1024 60.84 51.94 52.55 71.44 67.43
+SPQR-R 3.98 1024 58.11 49.96 50.54 68.49 63.46
+GWQ-O 4.63 1 63.12 54.08 55.94 75.33 70.36
+GWQ-R 3.98 1 60.78 51.48 52.46 71.43 66.54
Language Modeling Task. We first evaluate the perplexity and accuracy of the quantized model on the WikiText2 Merity et al. (2016) and C4 Raffel et al. (2020) datasets. As shown in Tab. 2 and Tab. 3, at the same average bits, the GWQ compressed model has lower PPL as well as higher Acc compared to the SPQR compressed model. Such results show that the quantized model with the same compression rate GWQ performs better in language modeling than the quantized model with SPQR.
When the average bit-width in the model after GWQ quantization is lower than 4, it still outperforms the 4-bit full-precision quantized model. This indicates that GWQ quantization is more robust compared to the full-precision approach.
For the Qwen-VL family Bai et al. (2023a, b) of models whose foundational model is not Llama, the quantized model of GWQ can still maintain strong language modeling ability. Thus, it is shown that GWQ can be generalized to different foundational models.
@@
Grounding Detection. The grounding detection tasks are mainly focused on zero-shot target detection. We choose the RefCOCO (Kazemzadeh et al., 2014; Yu et al., 2016) dataset to evaluate the zero-shot capability of the quantized model compared to other quantization methods. Intersection over Union $IoU$, which is denoted as:
@@
As shown in Tab. 3, Compared to SPQR, GWQ quantized same-precision Qwen-VL family models show a significant increase in detection accuracy. This indicates that the GWQ method is better suited for grounding detection scenarios compared to SPQR. Compared to the full-precision quantization method, the detection ability of the GWQ quantized model is still better than the models quantized by other methods when the average bits of the GWQ quantized model are less than 4. This suggests that the GWQ quantized model not only performs well on language modeling tasks but also has the potential to generalize to multimodal scenarios. The performance is shown in Fig. 3, for the task of visual grounding detection of long context, the GWQ post-quantization model is more advantageous than SPQR post-quantization.
Table 5: Evolution of vison-language question answering task datasets. GWQ outperforms other quantization methods on both LLaVA and Qwen-VL families. Results on ScienceQA dataset, Question classes: NAT is Natural Science, SOC is Social Science, LAN is Language Science, TXT is Text Context, IMG is Image context, NO = No context.
Accuracy (%)
Model Methods
Average Bit
Calib. Num
Average NAT SOC LAN TXT IMG NO
LlaVA-V1.5 Original 16 - 62.94 60.98 63.10 67.27 62.99 63.24 60.04
+OPTQ 4 1024 49.91 48.01 50.12 55.84 48.96 48.92 47.63
+QuIP 4 1024 51.76 49.98 51.92 57.78 50.69 50.63 49.58
+GPTQ 4 1024 52.69 51.01 52.95 58.64 51.73 51.71 50.15
+AWQ 4 512 54.66 53.97 55.37 61.36 54.70 54.66 52.29
+SPQR-O 4.63 1024 58.08 56.09 58.03 62.84 57.88 57.93 55.68
+SPQR-R 3.98 1024 55.23 53.99 55.36 60.36 54.68 54.67 52.31
+GWQ-O 4.63 1 60.55 58.89 60.68 64.98 60.24 60.56 57.94
+GWQ-R 3.98 1 56.99 55.01 57.10 61.82 56.78 56.84 54.36
Qwen-VL-chat Original 16 - 68.89 73.20 65.12 68.77 71.07 66.84 68.35
+OPTQ 4 1024 55.99 55.14 58.46 51.72 54.98 60.76 54.85
+QuIP 4 1024 57.75 56.98 60.02 53.58 56.82 62.34 56.76
+GPTQ 4 1024 58.65 57.99 60.64 54.25 57.84 63.24 57.96
+AWQ 4 512 61.39 60.93 63.32 57.15 60.55 65.90 60.46
+SPQR-O 4.63 1024 64.03 67.56 59.91 63.24 66.25 64.02 63.21
+SPQR-R 3.98 1024 61.38 60.92 63.33 57.14 60.56 65.91 60.43
+GWQ-O 4.63 1 67.39 70.16 62.20 65.91 68.72 66.14 65.85
+GWQ-R 3.98 1 64.03 67.66 59.98 62.12 68.33 63.96 62.14
Qwen-VL Original 16 - 66.70 63.55 66.48 69.02 67.91 66.25 66.99
+OPTQ 4 1024 54.03 50.89 52.97 57.03 56.69 52.98 53.64
+QuIP 4 1024 55.80 52.76 54.81 58.65 58.23 54.94 55.38
+GPTQ 4 1024 56.64 53.57 55.68 59.49 59.07 55.86 56.14
+AWQ 4 512 59.17 56.02 58.32 61.90 61.01 58.72 59.02
+SPQR-O 4.63 1024 61.81 58.64 61.93 64.14 62.43 61.76 61.98
+SPQR-R 3.98 1024 59.17 56.01 58.34 61.92 60.99 58.71 59.04
+GWQ-O 4.63 1 63.87 60.43 63.59 66.34 64.94 63.46 64.46
+GWQ-R 3.98 1 61.37 57.53 60.84 63.74 63.01 61.44 61.87
Table 6: efficiency comparison between different quantization methods. GWQ is not the best compared with other quantization methods.
Average Calibration Llama-2-7B-hf Llama-2-13B-hf Llama-3-8B-hf
bit number Speedup ($\uparrow$) Memory ($\downarrow$) Speedup ($\uparrow$) Memory ($\downarrow$) Speedup ($\uparrow$) Memory ($\downarrow$)
Original 16 - $\times$1.00 12.83G $\times$1.00 23.63G $\times$1.00 14.57G
+OPTQ 4 1024 $\times$1.90 3.21G $\times$1.91 6.9G $\times$1.90 3.84G
+QuIP 4 1024 $\times$1.98 3.11G $\times$1.99 6.71G $\times$1.98 3.61G
+GPTQ 4 1024 $\times$2.00 3.11G $\times$2.01 6.71G $\times$2.00 3.61G
+AWQ 4 512 $\times$2.01 3.10G $\times$2.00 6.72G $\times$2.01 3.61G
+SPQR-O 4.63 1024 $\times$1.28 N/A $\times$1.27 N/A $\times$1.28 N/A
+SPQR-R 3.98 1024 $\times$1.26 N/A $\times$1.25 N/A $\times$1.26 N/A
+GWQ-O 4.63 1 $\times$1.23 4.16G $\times$1.22 7.13G $\times$1.23 4.84G
+GWQ-R 3.98 1 $\times$1.22 4.16G $\times$1.21 7.13G $\times$1.22 4.84G
Massive Multitask Language Understanding. In this section, we will discuss the ability of Llama-2 and Llama-3 families using different quantization methods in massive multitask language understanding with MMLU dataset Hendrycks et al. (2021a, b).
As shown in Tab. 5, compared with SPQR, GWQ quantized Llama-2 and Llama-3 families show a substantial improvement in multimodal comprehension accuracy. This indicates that GWQ is more suitable for application in complex scenarios than SPQR. Compared with the full-precision quantization method, when the average bits of the GWQ quantized model are less than 4, the GWQ quantized model still outperforms the models quantized by the other methods in language understanding ability. The results show that our method maintains proficiency in single-task language modeling and also exhibits strong adaptability in complex multi-task scenarios.
Table 7: Effect of the number of calibration set to the quantization. Different numbers of calibration set affect a little to quantifies models’ perform. All calibration set are from Red Pajama1 dataset.
sample_number=1 sample_number=4 sample_number=16 sample_number=64 sample_number=512 sample_number=1024
[…]
Llama2-7B-hf 5.48 6.99 5.49 6.98 5.47 6.98 5.45 6.96 5.46 6.97 5.46 6.96
Llama2-13B-hf 4.90 6.48 4.91 6.47 4.90 6.47 4.89 6.46 4.89 6.48 4.89 6.48
Llama3-8B-hf 6.41 9.15 6.43 9.17 6.42 9.16 6.41 9.15 6.40 9.14 6.40 9.14
Qwen-VL 8.24 10.10 8.23 10.10 8.25 10.11 8.23 10.09 8.23 10.09 8.23 10.09
Qwen-VL-chat 9.86 11.88 9.86 11.88 9.85 11.86 9.85 11.87 9.85 11.87 9.85 11.86
Table 8: Effect of different calibration set to quantization results. different calibration set effects a little to quantization results, but special calibration set of task affects a lot.
WikiText2 C4 Red Pajama1 Red Pajama2 AWQ’s Calibration
[…]
Llama2-7B-hf 5.46 6.96 5.47 6.92 5.48 6.99 5.48 6.98 5.80 7.03
Llama2-13B-hf 4.85 6.47 4.88 6.85 4.90 6.48 4.91 6.47 5.99 6.51
Llama3-8B-hf 6.38 9.14 6.40 9.10 6.41 9.15 6.41 9.16 6.49 9.56
Qwen-VL 8.22 10.08 8.23 10.05 8.24 10.10 8.23 10.11 8.31 10.17
Qwen-VL-chat 9.84 11.89 9.86 11.86 9.86 11.88 9.87 11.87 9.93 11.94
Vision-Language Question Answering. The vision-language question answering (VQA) tasks aim to assess the model’s ability to generate accurate responses to questions based on visual and textual inputs. We evaluate the performance of GWQ and other quantitative methods applied to LlaVA-V1.5 Liu et al. (2024), Qwen-VL Bai et al. (2023a), and Qwen-VL-chat Bai et al. (2023a) models on the ScienceQA dataset Lu et al. (2022).
As shown in Table 4, similar to the previous task, the performance of the GWQ quantized model achieved SOTA on VQA accuracy, compared to the other methods. This suggests that GWQ is highly effective for vision-language question-answering tasks. Even when the average bits of the GWQ quantized models are below 4, the GWQ quantized models still outperform models quantized by other methods. This indicates that the GWQ quantized models not only maintain strong performance in language modeling but also generalize well to complex multimodal tasks like VQA.
Model Efficiency Comparison. In this section, we will discuss the impact of different quantization methods on model acceleration and compression effects.
@@
For the model throughput (tokens/s), the speedup rate can be obtained by calculating the time required to generate 512 tokens after quantization and comparing it with the original model. For the model inference overhead (G), we use the peak memory when the model generates 512 tokens as the maximum inference memory by recording it. As shown in Tab. 6, the model after quantization using GWQ 4bit achieves about 1.2$\times$ of speedup compared to the original model. However, compared to SPQR, the acceleration effect of GWQ is not significant. We believe that the insignificant effect of GWQ is due to the fact that the outlier values saved by GWQ search are sparse, which is not favorable for inference calculation. Compared with the full-precision quantization methods, GWQ and SPQR are not advantageous due to their hardware adaptability.
@@ 4.3 Ablation Study
In this section, we test the number of checksets and the effect of different checksets on our quantization method and justify our choice of parameters.
Effect on Number of Calibration. We select different number of samples from the Red Pajama dataset as calibration samples respectively, by taking the gradient mean and searching for the corresponding outlier of the model.
We tested the performance of the model after quantizing different numbers of samples on the WikiText2 and C4 datasets. As shown in Tab. 7, we find that the model effect fluctuates slightly when the number of calibration samples is small, but it does not affect the final quantitative results very much. At the same time, when the number of calibration samples is increased, we find that the effect of the quantized model is similar to that when there is only one sample. Therefore, from the consideration of data efficiency and method robustness, we believe that it is optimal to choose a sample of 1.
Effect on Different Calibration. We choose different calibration data Lin et al. (2024a); Weber et al. (2024); Merity et al. (2016); Raffel et al. (2020) to calibrate the Llama families model Dubey et al. (2024); Touvron et al. (2023b) and test the PPL of the quantized model on the WikiText Merity et al. (2016) and C4 Raffel et al. (2020) datasets.
As shown in Tab. 8, different types of calibration samples do not have a significant effect on the quantized model. We guess that the impact of the calibration data on the quantization is less due to the fact that when collecting the gradients, we used the gradient mean as the outlier indexing basis. However, for specialized calibration sets, such as using the C4 calibration set for quantization calibration, and when testing on the C4 calibration set, the model will show better results. At the same time, we believe that the quantization effect is related to the complexity of the checksum set. When WikiText is used as the check set, its quantized model effect has a slight decrease compared to the rest of the models. This may be due to the fact that the language structure of the WikiText check set is relatively simple, which cannot realize more accurate calibration.